The following account balances were included in Bromley Company’s balance sheet on December 31, 2018:

Land $100,000
Land improvements 20,000
Buildings 300,000
Machinery and Equipment 500,000

During 2019, the following transactions occurred:

Jan. 1 Land was acquired for $70,000 for a future building site. Commissions of $4,000 were paid to a real estate agent.
Feb. 1 A factory and land were acquired from Kent Development Company by issuing 20,000 shares of $3 par common stock. At that time, the stock was selling for $10 per share on the New York Stock Exchange. The independently appraised values of the land and the factory were $60,000 and $180,000, respectively.
Mar. 1 Equipment was acquired at a cost of $120,000. In addition, sales tax, freight costs, and installation costs were $7,000, $10,000, and $16,000, respectively. During installation, the equipment was damaged, and $2,000 was spent for repairs.
Apr. 1 A new parking lot was installed at a cost of $30,000.
Aug. 1 Half the land purchased in Item 1 was prepared as a building site. Costs of $26,000 were incurred to clear the land, and the timber recovered was sold for $3,000. A new building was built for $60,000. Architect’s fees relating to construction were $18,000, and imputed interest on equity funds used during construction was $15,000. No debt is outstanding.
Sept. 1 Costs of $20,000 were incurred to improve some leased office space. The lease will terminate in 2021 and is not expected to be renewed.
Oct. 1 A group of new machines was purchased under a royalty agreement that provides for payment of annual royalties based on units produced. The invoice price of the machines was $30,000, freight costs were $2,000.
Dec. 31 Royalty payments on the new machines for 2019 totalled $12,000.
Required:

Prepare journal entries to record all the preceding events. Unless otherwise indicated, assume the company makes all payments in cash

Prepare journal entries to record all the events that occurred during 2019. Unless otherwise indicated, assume the company makes all payments in cash.

General Journal Instructions

PAGE 16

GENERAL JOURNAL

DATE ACCOUNT TITLE POST. REF. DEBIT CREDIT

1

2

3

4

5

6

7

8

9

10

11

12

13

14

15

16

17

18

19

20

21
Answer #1

Answer:

No. DATE ACCOUNT TITLE POST. REF. DEBIT CREDIT
1 Jan-01 Investment in Land 74,000
Cash 74,000
To record the land acquired
2 Feb-01 Land 50,000
Buildings 150,000
Common stock 60,000
Additional Paid-In capital - Common Stock 140,000
To Record the buildings land acquired through issuance of common stock
3 Mar-01 Equipment 153,000
Repair and Maintenance Expense 2,000
Cash 155,000
To record expense associated with equipment
4 Apr-01 Land Improvements 30,000
Cash 30,000
To record the instalment of new parking lot
5 Aug-01 Land 60,000
Buildings 78,000
Investment in Land 37,000
Cash 101,000
To record the costs associated with land and buildings
6 Sep-01 Leasehold Improvements 20,000
Cash 20,000
To record the Leasehold Improvements
7 Oct-01 Equipment 32,000
Cash 32,000
To record the purchase of equipment
8 Dec-01 Royalty Expense 12,000
Cash 12,000
To record the royalty payment

Calculation:

Entry #1

Land Acquired and cash paid = Cost of acquiring + Commission = $70,000 + $4,000 =74,000

Entry #2

Amount of Land = (Appraised value of land/Total appraised value of land and building)*Common stock = (60000/(60000+180000))*20,000*10 = 50,000

Amount of Building = (Appraised value of Building/Total appraised value of land and building)*Common stock = (180000/(60000+180000))*20,000*10 = 1,50,000

Common stock = 20000*$3 = 60,000

Entry #3

Value of Equipment = Cost of acquiring +sales tax + freight costs + installation cost=120000+7000+10000+16000 = 1,53,000

Entry #5

Costs associated with land = Half of cost of land acquired + cost to clear the land + half of commission relating to land - sale of timber recovered = (35000+26000+2000)-3000 = 60,000

Costs associated with land = Cost of construction + architecture fee = 60000+18000 = 78,000.

Investment in Land reduced (half of land acquired)= 74000/2 = 37,000